Some recent financial statements for Smolira Golf, Inc., follow.

SMOLIRA GOLF, INC.
Balance Sheets as of December 31, 2015 and 2016
2015 2016 2015 2016
Assets Liabilities and Owners’ Equity
Current assets Current liabilities
Cash $ 3,061 $ 3,057 Accounts payable $ 2,178 $ 2,650
Accounts receivable 4,742 5,731 Notes payable 1,775 2,166
Inventory 12,578 13,732 Other 95 112
Total $ 20,381 $ 22,520 Total $ 4,048 $ 4,928
Long-term debt $ 12,900 $ 15,660
Owners’ equity
Common stock and paid-in surplus $ 40,500 $ 40,500
Fixed assets Accumulated retained earnings 15,679 39,477
Net plant and equipment $ 52,746 $ 78,045 Total $ 56,179 $ 79,977
Total assets $ 73,127 $ 100,565 Total liabilities and owners’ equity $ 73,127 $ 100,565
SMOLIRA GOLF, INC.
2016 Income Statement
Sales $ 188,370
Cost of goods sold 126,703
Depreciation 5,283
EBIT $ 56,384
Interest paid 1,380
Taxable income $ 55,004
Taxes 19,251
Net income $ 35,753
Dividends $ 11,955
Retained earnings 23,798

Smolira Golf has 11,000 shares of common stock outstanding, and the market price for a share of stock at the end of 2016 was $82.
  
What is the price–earnings ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Price–earnings ratio             times
  
What is the price–sales ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Price–sales ratio             times
  
What are the dividends per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Dividends per share            $
  
What is the market-to-book ratio at the end of 2016? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Market-to-book ratio
Answer #1
Price–earnings ratio Market price per share/Earnings per share $ 25.23 =82/(35753/11000)
Price–sales ratio Market price per share/Sales per share $    4.79 =82/(188370/11000)
Dividends per share Dividends paid/Number of shares $    1.09 =11955/11000
Market-to-book ratio Market price per share/Book value per share $ 11.28 =82/(79977/11000)